OpenStudy (anonymous):

If c + 12 = 37, what is the value of c?

OpenStudy (alexandervonhumboldt2):

subtract 12 from both sides: c+12-12=37-12

OpenStudy (mynameisnemo):

To solve this, you need to get c alone by subtracting 12 from both sides.

OpenStudy (mynameisnemo):

c + 12 = 37 - 12 - 12 c= 25